OpenStudy (anonymous):

Find the ratio of areas of two circles if the ratio of their radii is 3:4.

OpenStudy (anonymous):

would it be 3:4 only?

OpenStudy (zepp):

It should be 3^2:4^2

OpenStudy (lgbasallote):

hmm

OpenStudy (zepp):

If the ratio of length is \(k\) then ratio of area is \(k^2\) then ratio of volume is \(k^3\)

OpenStudy (anonymous):

9:16 ?

OpenStudy (anonymous):

Let r1=3k and r2=4k then area of circle of r1 = pi*(3k)^2 = pi*9k^2 area of circle of r2 = pi*16k^2 then ration of circle of r1 and r2 = pi*9k^2 : pi*16k^2 = 9:16

OpenStudy (zepp):

In this case \(\large k=\frac{3}{4}\) The area would be \(\large k^2=(\frac{3}{4})^2=\frac{9}{16}\) Exactly :)

OpenStudy (anonymous):

Thanx both of you :)

OpenStudy (zepp):

You are welcome :)